OpenStudy (anonymous):

number between 1 and 5/6

OpenStudy (solomonzelman):

\(\Large\color{red}{ \bf 1=\frac{12}{12} }\) \(\Large\color{red}{ \bf \frac{5}{6}=\frac{10}{12} }\) need more help ?

OpenStudy (anonymous):

These are the options I can choose from a 6/7,b 4/5,c4/7 ,d11/6